### v1

# Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions

Abstract
Urban traffic congestion, worsened by the rapid urbanization and the increasing prevalence of private vehicles, has significantly increased commuting time for everyone. In this paper, we used a dataset with over 400,000 real mobility trajectories of individuals spanning 9 days in a major Chinese city to investigate an innovative approach to swap homes between households in addressing the challenge of peak-hour traffic congestion. We observed that, empirically, households choose their home location strategically such that the average commuting distance is roughly $3$ times less than that when their home is randomly located, showing features of self-organization. Remarkably, we found that the average commuting distance can be further reduced by $50\%$ through home swapping at the city-level, leading to a large reduction in traffic congestion. To make home swapping more realistic, we swap homes only if the following socio-demographic factors including the distance from the city center, housing price and amenity accessibility are preserved for both households, such that the average commuting distance can still be reduced by $13\%$. As both home-workplace distance and traffic congestion are reduced, as a side benefit, carbon emissions from vehicles are also greatly reduced by almost $80\%$, and $40\%$ when socio-demographic factors are considered. The distance from the city center is shown to be the most influential factor affecting the benefit brought by home swapping, and further analysis indicates that developing a polycentric city layout could significantly enhance such benefit. This study suggests that mitigating traffic congestion requires a long-term, holistic and strategic approach to urban planning, suggesting a need for coordinating individual residence locations and a polycentric city layout.
I Introduction
With the rapid urbanization, an increasing ratio of the world population resides in urban areas sun2020dramatic . While the room for expansion in urban road infrastructure is small to cope with the increasing population, traffic congestion becomes more severe ccolak2016understanding ; pi2019visual ; zhou2022identifying ; dasgupta2021spatiotemporal ; wang2023aggravated ; manfredi2018mobility ; dahlmanns2023optimizing ; zhang2019scale . This does not only reduce the daily comfort of city commuters, but also waste a lot of resources in terms of time and energy, and induce air and noise pollution, impacting the socio-economy of cities at a holistic level.
Specifically, traffic congestion can be divided into non-recurring and recurring congestion. Non-recurring congestion is caused by unexpected events and is usually unpredictable. Recurring congestion, on the other hand, is caused by periodic traffic flows, such as the commuter flow during morning and evening peak hours. This type of congestion is regular and predictable and may constitute a major part of the urban congestion afrin2020survey . To alleviate recurring traffic congestion, many cities have developed a series of measures. These include advocating the use of public transport, staggered work hours, establishing road tolls in peak hours, restricting vehicle usage based on license plate numbers, etc. In addition, researchers have proposed various methods to ease traffic congestion, including traffic flow prediction  ren2014predicting ; hou2019adaptive ; ma2021short , control of traffic lights  cao2016unified ; de2020development , time-based travel arrangements  xu2018planning ; yildirimoglu2021staggered and dynamic traffic assignment  aghamohammadi2020dynamic ; yeung2013physics ; yeung2019coordinating . However, most methods are remedial without fundamentally tackling the source of traffic congestion.
As we all know, the root cause of traffic congestion during peak hours lies in the need of individuals to travel between home and their workplace or school  liu2018investigating . The longer the commuting distance, the more likely private motor vehicles are used, which leads to more vehicles on roads, and hence increased traffic congestion during peak hours  wheaton2004commuting ; zhao2020long . On the contrary, if the commuting distance is short, people can use non-motor vehicles such as bicycles or even walk to their workplace. Thus, there were studies which examine residences, workplaces, and commuting distances using empirical datasets  huang2018tracking ; plaut2006intra ; gerber2020workplace ; surprenant2013commuting ; guo2021fractal ; suzuki2012jobs ; zhao2011impact ; asmussen2024interplay ; hu2017different ; ta2017understanding ; shen2021job . Studies show that moving or changing jobs can change the spatial arrangement of residences and workplaces for single-commuter and dual-commuter households, reducing their commuting distance and time  huang2018tracking ; surprenant2013commuting . However, these studies were usually based on surveys, which have a finite number of respondents and hence do not fully reveal the benefit brought by reallocation of residence at the city level.
In this paper, we examine the impact of city-wide home-swapping in reducing traffic congestion and alleviating carbon emission. We used more than 400,000 real individual mobility trajectories over 9 days from a Chinese city, to infer the home and workplace location as well as household members of each individual, and identify all the commuter path during peak hours using Baidu map BaiduPlatform . We found that randomly re-shuffling residence locations increases the average commuting distance by $2.83$ times, indicating a strong self-organization in the actual choices of residence location. By strategic home swapping, urban congestion during peak hours is greatly reduced since the average commuting distance is reduced by $50.4\%$, which complies with the notion of a 15-minute city abbiasov202415 ; khavarian202315 ; papadopoulos2023measuring . Our study is the first to verify the benefit brought by home-swapping using such a real and big data.
To make home swapping more realistic, we swap home only if the following three socio-demographic factors are preserved for both households, namely the distance from the city center, housing price and amenity accessibility. In this case, the average commuting distance can still be reduced by 13%. Owing to the combined effect of shorter commuting distance and reduced traffic congestion, carbon emissions from vehicles are also greatly reduced by $78\%$, and by $43\%$ when socio-demographic factors are considered. Among these, the distance from the city center influences home-swapping most, followed by the amenity accessibility and housing price. Based on our findings, we found that establishing a polycentric, i.e. multi-center, city layout can effectively increase the benefit brought by home swapping.
II Results
Our big data of empirical human trajectory are based on the 4G communication records between base stations and mobile phones provided by one of the three major service providers in Shijiazhuang, the capital of Hebei province in China, from May $22^{\rm nd}$ to $28^{\rm th}$, 2017, and from June $3^{\rm rd}$ to $4^{\rm th}$. The original data include records from around 3 million users out of a total population of 10 million in the city and spans 9 days, including 2 weekends and 5 weekdays. To identify meaningful locations in the user trajectories, we divide the whole period covered by the dataset into 15-minute intervals (see Materials and Methods). If an individual stayed in the same location for more than 10 minutes in each interval, the location was recorded as a stop for that user. In addition, when we compute the similarity between the trajectories of individuals later, the division into time intervals facilitates the comparison of defined stops of individuals in each time interval. Other than this high-frequency mobility dataset in Shijiazhuang, we also used the dataset of residence and workplace from Shenzhen of China and the dataset of call detail records (CDR) from Abidjan of Cote d’Ivoire to cross-validate the results observed in Shijiazhuang (see Section S1 of the Supplementary Information (SI) for details).
To study the benefit brought by home swapping, one has to obtain two pieces of essential information for each individual, namely (1) his/her location of residence and workplace, and (2) the members in his/her household who also commute. First, we used an individual’s trajectory data from the 5 weekdays to infer his/her residence and workplace locations (Fig. 1A). In the subsequent analyses, for each individual, we infer his/her (1) residence location, based on the location he/she stayed for at least four nights in the 5 weekdays, (2) workplace location, based on the location where he/she stayed for at least 2 hours each day and spends the most time during the daytime in the 5 weekdays. Given that traffic congestion in urban central areas is particularly severe, we only studied the main urban area in Shijiazhuang, which contains nearly 420,000 anonymous commuter trajectory records (see Materials and Methods). Although individuals’ residence and workplace locations are only inferred, such inference based on cell phone locations should be accurate; this leads to a much larger dataset compared to the data analyzed in conventional studies which are mainly based on surveys.
Next, we infer the household membership between individuals. Since (1) members in the same household usually have different workplace locations but they may enjoy the weekend outing together, and (2) previous studies demonstrated that the intimacy between individuals is related to the similarity between their mobility trajectories  wang2011human , we use the similarity between the trajectories of two individuals with the same residence location from the 4 days in the 2 weekends to evaluate their household membership. Specifically, we denote the similarity between the mobility trajectory of individual $i$ and $j$ in a single day by $S_{ij}$ (see Materials and Methods). The closer the value of $S_{ij}$ to $1$, the more similar their trajectory, and closer the relationship between $i$ and $j$. Nevertheless, even in weekends, family members may only spend part of the day together rather than the entire day. Therefore, we determine if $i$ and $j$ are household members by using the highest $S_{ij}$ among the four values found in each of the 4 days of the 2 weekends. If $S_{ij}>0.5$, individual $i$ and $j$ are considered members in the same household. A schematic example of computing the similarity between the trajectories of two individuals is shown in Fig. 1A.
To make our inference more realistic, we ensure the inferred distribution of household size agrees with that found by the census in Shijiazhuang, by grouping individuals $i$ and $j$ with no household members but with higher $S_{ij}$ to form households, more details can be found in Sec. S2 of the SI. The commuting routes between their home and the workplaces of three members of an inferred household are shown in Fig. 1B.
To start analyzing the potential benefit brought by home swapping, we first show the spatial distribution of the location of residence and workplace of commuters in the city in Fig. 2A and Fig. 2B respectively. We divide the urban area of Shijiazhuang into a grid of squares with a dimension of $500m\times 500m$; the more the residence or workplace locations in a grid, the higher the intensity of the color on the heat maps in Figs. 2A and 2B. As we can see, areas close to the city center (e.g. Yuhua ,Chang’an ,Xinhua and Qiaoxi Districts) have a high concentration of residence and workplace locations, and the opposite is observed in the surrounding suburban areas. Residence locations are more dispersed and numerous, while workplaces are more concentrated near the city center (see Fig. 2C). The KL divergence (see Materials and Methods) between the two spatial distributions is 0.096, indicating that they are very similar. Similar observations are obtained in Shenzhen and Abidjan and are shown in Fig. S16 and S20 respectively.
Despite the similarity between the overall spatial distributions of residence and workplace locations, it does not necessarily imply that commuters live near their workplaces. As shown in Fig. 2D, the distribution of their daily commuting distance is roughly exponential, such that many commuters commute a long distance daily, suggesting a significant room for exploitation. To examine whether commuters used motor vehicles for commuting, we found that the travel distance of Shijiazhuang residents on-demand vehicles peaked at 2.5 km (see in Fig. S1). We therefore assume that 2.5 km is the shortest distance for commuters to use motor vehicles for commuting. In addition, we also shown in Sec S6 of the SI a different threshold does not qualitatively affect the obtained results. Based on the empirical home-workplace distance, almost $64\%$ of the Shijiazhuang citizens commute a distance larger than 2.5 km and thus they are assumed to use motor vehicles for commuting. In addition, we see that the inferred departure time of commuters from their home peaks within a short period during the morning rush hours in Fig. S2, and together with the long commuting distance, are the dominant factor for urban traffic congestion according to  wheaton2004commuting ; yildirimoglu2021staggered .
One major reason underlying the long commuting distance for some commuters, is that they live in a location where their other household members have a short commuting distance. We further analyze how households with multiple members choose their home location. In Fig. 2E, we showed the distribution of the difference $d_{\rm max}-d_{\rm min}$ between the largest and the smallest commuting distance among members in the same household. As we can see, the number of households generally decreases with $d_{\rm max}-d_{\rm min}$, with more households found at the small $d_{\rm max}-d_{\rm min}$. The highest peak at the smallest $d_{\rm max}-d_{\rm min}$ implies that many households set their residence location at the centroid among the workplace locations of all the household members, such that $d_{\rm max}-d_{\rm min}$ is minimal.
Interestingly, we note a second peak at an intermediate $d_{\rm max}-d_{\rm min}$, implying that many other households reside at a location closer to one of the household members. As shown in Fig. 2F, dual-commuter households exhibit a clear bifurcation, i.e. some households choose to live in between workplaces such that $(d_{\rm max},d_{\rm min})$’s lie near the diagonal, while others opt for locations close to one member but farther away from the other with $(d_{\rm max},d_{\rm min})$’s lying near the horizontal axis. With the increase in the number of household members, it becomes less feasible for families to live close to everyone’s workplaces. In this case, we see that $(d_{\rm max},d_{\rm min})$’s are mainly found near the horizontal axis, implying that they are more inclined to reside in locations close to the workplaces of a few members, sacrificing at least one member’s commuting time, consistent with the conclusion in  surprenant2013commuting .
Home Swapping. Given the household membership and members’ workplaces remain unchanged, we examine the effectiveness of an innovative approach to swap homes among households to reduce the total commuting distance in the whole city. To anticipate the potential benefit brought by such an approach, as shown in Fig. 2A-C, the spatial distributions of home and workplace locations are very similar in the urban area, but most commuters’ home and workplace are not in the same location. This means that their home location is someone else’s workplace, and their workplace location is someone else’s home. As we show in Fig. 2F, it is impossible to find a home location close to every household member’s workplace, nevertheless, we still anticipate that after strategic home swapping, the sum of the commuting distance for the swapped households can be reduced, leading to an overall reduction in the average commuting distance at the city level.
Random Home Swapping (RHS). As a benchmark, we randomly swap homes among households and examine its impact on the commuting distance of the city. The results in Fig. 3A shows that the average commuting distance after random home swapping is $2.83$ times higher than that computed based on the real data. It is worth noting that the commuting distance for each commuter in the real data was obtained through crawling the actual commuting path on Baidu Map (see Materials and Methods). This indicates that real households do exhibit an extent of self-organization in choosing home locations, taking into account the commuting distance to each household member’s workplaces. Other factors such as transportation, income, amenities around the residence, etc. may also be considered. However, we expect that there is still a large room for improvement at the holistic city level as the empirical data only represent self-organization at the household level.
Greedy Home Swapping (GHS). To examine the benefit brought by strategic home swapping, we propose a greedy strategy to swap homes only if the total commuting distance of the two swapped households decreases. Due to the restriction in living space, we only allow home-swapping between 1-commuter and 2-commuter households, as well as swapping between 3-commuter and 4-commuter households. We terminate the swapping attempts if no households can be swapped to reduce commuting distance for a consecutive $1\times 10^{6}$ attempts. As the goal of home-swapping is effectively to match each household to a residence location given the workplace locations of all citizens, the Hungarian algorithm is able to identify the optimal solution for this matching problem and hence the optimal allocation of residence locations for all household. Nevertheless, its computation complexity of $O(N^{3})$ and the large number of $2.4\times 10^{5}$ households in our real data would make the matching intractable. Therefore, we employed the above swapping procedure to reduce the commuting distance of the swapped households and the city as whole.
In this case, after home swapping, the decrease in commuting distance fundamentally reduces traffic congestion, through the combined effect of a shorter traveling distance and a reduced number of motor vehicles used since some commuters may walk or cycle instead. An example of the impact of greedy home swapping (GHS) on individual households is illustrated in Fig. 1C-D. In Fig. 3A, we further show the average commuting distance over all commuters in the city based on the real data, after random and greedy home swapping, which we denote by $\bar{d}_{\rm real}$, $\bar{d}_{\rm random}$ and $\bar{d}_{\rm greedy}$ respectively. As we discussed before, $\bar{d}_{\rm random}=2.83\bar{d}_{\rm real}$, and after greedy home swapping, $\bar{d}_{\rm greedy}=0.5\bar{d}_{\rm real}$. To better compare the benefits brought by different home swapping approaches later, we denote $\Delta d$ as the percentage decrease in the average commuting distance over all commuters, given by
$\Delta d=\frac{\bar{d}_{\rm real}-\bar{d}}{\bar{d}_{\rm real}},$ (1)
where $\bar{d}_{\rm real}$ and the $\bar{d}$ are the original and the new average commuting distance over all urban commuters before and after home swapping, respectively. After GHS, $\Delta d=50.36\%$, which is a highly significant reduction.
In Fig. 3B, we further show the scatter plots of the average commuting distances in a household before and after GHS. We find that the average commuting distance for most households are shortened. Nevertheless, some households with already a short commuting distance have their commuting distance increased after GHS, but remains within an acceptable range. It is because the households which swap home with these families have a larger reduction in commuting distances compared to their increases. This is also an example of individual optimum versus social optimum, such that some individual households have to sacrifice to choose a sub-optimal home location to benefit others. Finally, there are also a few households whose commuting distances have not changed, either because they did not swap home or the commuting distance between their new home and workplaces remain unchanged.
After households’ residences are re-allocated, there are changes in their routes, potentially their modes of transportation, and departure times (see Materials and Methods). For instance, commuters with a commuting distance of less than 2.5 km may switch to a non-motorized travel such as bicycles. In this case, the congestion on different road segments at any specific time of the day may change.
To quantify the change in congestion, we define a city-level congestion index, called the average congestion coefficient $I_{t}$ (see Materials and Methods). We then analyze this index at different times of the day based on the real dataset, and after random and greedy home swapping. In Fig. 3C, for all periods of time, we found that the average congestion coefficient after RHS is highest, while that after GHS is lowest, with the empirical one in between. Further analysis was conducted on the traffic condition during the morning peak hours of 7-8 AM, when most commuters leave home for work. In the case of RHS, most road segments are severely congested and vehicles move slowly. After GHS, only a few road segments remain congested as shown in Fig. 3D-F; based on these results, one can anticipate that GHS alleviate congestion throughout the rush hour (see Fig. S4).
To verify the above remarkable benefit brought by GHS in cities other than Shijiazhuang, we conducted similar simulations using datasets obtained from Shenzhen and Abidjan and obtained qualitatively similar results and benefits as in Fig. 3A (see SI section 7 for details). These results show that there are great benefits brought by GHS in the cities examined, which can be generic among all other cities.
Greedy Home Swapping with Socio-demographic Consideration (SDGHS). While Greedy Home Swapping is based solely on reducing commuting distance which significantly alleviates traffic congestion, previous literature has confirmed that socio-demographic factors play a vital role in people’s choice of residence  tillema2010influence ; guo2020modeling ; rouwendal2001preferences ; ho2016workplace ; hu2017different ; guan2020multiplicity ; liu2022quantifying ; wang2024does ; al2024pathway ; cao2017examining ; blumenberg2023housing . Families, although not fond of spending time in transit, have preference for specific housing characteristics that are strong enough for them to accept a longer commuting time. In view of this socio-demographic consideration, to make home swapping more realistic, we further consider three factors during home swapping: (1) distance from the city center, (2) housing price, and (3) amenity accessibility.
Specifically, on top of reducing commuting distance, two households swap their home only if
  1. 1.
    the difference in the distance between their home and the city center is less than $T_{\rm center}$ km (c-GHS) - It is because the distance from the city center is an important factor when families choose their home. According to ref. xu2023urban and the data, we define the most visited location in Shijiazhuang to be its city center, i.e. the largest central business distract , i.e. the Beiguo Shopping Mall in Shijiazhuang.
  2. 2.
    the difference between the housing prices of their home is less than $T_{\rm price}\%$ (p-GHS) - Since housing prices affect their choice of residence, for instance, they generally choose residence within their financial capacity. Here, we use the housing prices of those second-hand properties to represent housing price in that region (see Materials and Methods).
  3. 3.
    the difference between the availability of amenities around the two residences is less than $T_{\rm amenity}$ (a-GHS) - Since households consider the surrounding living environment when they choose their home abbiasov202415 , here we use the information of the Points of Interest (POI) around the corresponding residence to represent its living environment. For locations without the information of their surrounding POI (e.g. mostly villages), we do not swap their home with others (see Materials and Methods).
Figures 4A-C show the percentage decrease in the average commuting distance, i.e. $\Delta d$, after Greedy Home Swapping by consider each of above three socio-demographic factors, denoted as c-GHS, p-GHS and a-GHS respectively. We find that as the households’ tolerance on the three factors increases, $\Delta d$ also increases, which is reasonable as the socio-demographic consideration limit the flexibility in home swapping.
In Fig. 4D, we further examine $\Delta d$ under the joint consideration of all the three socio-demographic factors with different level of tolerance, which we denote by Socio-demographic Greedy Home Swapping (SDGHS). As we can see, $\Delta d$ reaches $25\%$ under high tolerance, i.e. $T_{\rm center}=$ 1000 meters, $T_{\rm price}=25\%$ and $T_{\rm amenity}=0.5$, but achieves only $3\%$ under low tolerance, i.e. $T_{\rm center}=$ 100 meters, $T_{\rm price}=5\%$ and $T_{\rm amenity}=0.1$. In Fig. 4E, we compare the average congestion coefficient $I_{t}$ computed based on the inferred residence and workplace location from the real dataset and the path of each commuter recommended in Baidu Map during morning and evening rush hour, with those after home swapping under the forementioned socio-demographic consideration with different tolerances. We found that the status of $I_{t}$ under strict consideration (low tolerance) of the factors is closer to that in the real data, while the congestion under loose consideration (high tolerance) has improved as shown in Fig. 4E. $I_{t}$ under medium consideration (medium tolerance) in rush hours is shown in Fig. S5. We also conducted similar tests using datasets obtained from Shenzhen and Abidjan, and obtained qualitatively similar results (see Fig. S18, S19A and S23A).
Finally, in addition to the average congestion coefficient, we analyze the degree of congestion on each road after GHS taking different factors into account, to provide a clearer understanding on how traffic congestion is relieved. As shown in Fig. S6, home swapping can effectively reduce the number of congested and slow-moving road segments during the rush hour. In addition, we show in Fig. 4F that there is a larger reduction of road congestion between the empirical congestion and that after home swapping under the consideration of housing price and amenity accessibility with medium tolerance. However, after GHS taking into account the distance to the city center, some roads remain similarly congested, implying that the distance to the city center is the main consideration of households which hinder the reduction of city-wide commuting distance by GHS. Further comparison among the reductions in average commuting distance based on any 2 out of the 3 socio-demographic factors is shown in Fig. S7, and the results also suggest $T_{\rm center}$ emerges as the major factor influencing GHS to improving urban congestion.
Finally, we examine one important assumption of GHS, such that all households are willing to swap their home as long as the total commuting distance of the two swapping households decreases, even at the expense of an increase in the commuting distance of one of them. Here, we assume that the households with an increase of commuting distance have an $x\%$ willingness to swap home but $1-x\%$ to reject. As shown in Fig. S8 of the SI, there is no significant impact on $\Delta d$ as willingness $x$ changes.
Reduction in Carbon Emissions. As a by-product, Greedy Home Swapping which reduces the average commuting distance and hence the traffic congestion, also reduces carbon emissions. In general, the longer the commuting distance, the longer the time vehicles stayed on the roads, the more congested were the roads, and the more emissions were produced  kissinger2019detailed ; xue2022impact ; muniz2018urban ; tian2024spatial ; bohm2022gross ; li2023assessing . To reveal the impact of home swapping on reducing carbon emissions, we examine the total carbon emissions (see Materials and Methods) in various cases as shown in Fig. 5A. As we can see, the total carbon emissions in RHS were much larger than the total empirical carbon emissions, and the total carbon emission in GHS was the least, such that the reduction in carbon emissions reaches $77.3\%$. We further compare road carbon emissions under RHS, real data, and GHS in rush hour in Fig. S9.
For GHS with socio-demographic consideration, we see in Fig. 5A that carbon emissions are significantly higher, such that the total emissions are higher in c-GHS than those of p-GHS or a-GHS. The total emissions are similar in the case of p-GHS and a-GHS, and the total emissions after SDGHS under medium tolerance are much higher than that with a single constraint. As shown in the upper panel of Fig. 5C, after SDGHS with medium tolerance, specifically $T_{\rm center}$ = 0.5km, $T_{\rm price}$ = 15% and $T_{\rm amenity}$ = 0.3, the carbon emissions of some roads in the city where have decreased significantly compared to the emissions estimated by the original un-swapped data. When we consider only the constraints of $T_{\rm price}$ and $T_{\rm amenity}$, the lower panel of Fig. 5C shows much more roads with reduced carbon emissions. By comparing the insets of the upper and lower panels, we see that the reduction of emissions of the corresponding roads decrease significantly when the distance to the city center is considered in home-swapping in the upper panel, in line with the conclusion in Fig. 4F, implying that the distance to the city center plays a large role on GHS in reducing total emissions. Finally, we show the reduction in carbon emissions during the morning and the evening peak hour in each case in Fig. S10, such that similar results are obtained.
Next, we analyze the dependence of the reduction in carbon emissions on the reduction in the average commuting distance, i.e. $\Delta d$, after GHS, as shown in Fig. 5B. As $\Delta d$ increases, the emission reduction continuously increases, with GHS reducing $30\%$ of the average commuting distance, leading to a $61\%$ reduction of carbon emissions. The nonlinear relation between the reduction in carbon emissions and the reduction in commuting distance has several origins. First, the changes in commuters’ departure time and route after home swapping change the congestion status of individual road segments. The reduction in congestion decreases the commuting time for a given commuting distance, thereby reduces carbon emissions. Additionally, as we have discussed before, when commuters change their residence locations, their type of travel also changes. When the commuting distance is less than 2.5 km, commuters may switch from driving to walking, leading to zero carbon emission for this group. We show the reduction of carbon emissions of individuals after GHS in Fig. S11, suggesting a potential trifurcation. Therefore, shortening average commuting distance does not only bring convenience to the residents’ daily lives but also hold a strong significance for protecting the urban environment and thus benefiting the health of the urban population.
Exploiting GHS by Polycentric Urban Structure. As we know, polycentric city structure affects urban traffic nilforoshan2023human ; jun2020effects ; zuo2024interactions ; rong2022inspection ; louail2014mobile ; louf2013modeling ; pan2024whither ; sun2020effects ; zhao2024unravelling . The above analysis also suggests that the distance to the city center is the primary factor affecting the choice of the residence location of households and thus the reduction of commuting distance by home swapping. We thus study whether a polycentric urban structure could reduce significantly the average commuting distance. In this case, by shifting from a monocentric city model, where most activities are concentrated in a single region, to a polycentric model, with multiple centers of activity across the city. This approach could lead to a more balanced spatial distribution of living-working locations, allowing a more efficient use of transportation infrastructure, reducing travel times and carbon emissions, improving accessibility to amenities and services, and hence enhancing the quality of life for urban residents.
In this analysis, several new activity centers are added to Shijiazhuang in addition to the original city center, each in a relatively bustling area of the city, with a distance of more than 3 km between each pair of centers. We assume that this is a policy-driven scenario, since city centers are planned by the city government at specific locations such that the development of infrastructures will follow, and residents are attracted to these new city centers owing to the expected development.
Specifically, when SDGHS was conducted with medium tolerance, we found that $\Delta d$ increases with the number of city centers. To examine the impact of the distance between the city centers, we study the scenarios when the new centers are the most visited locations outward from the original city center at a minimum distance of 3km, 6km, 9km, and 12km, respectively, as shown in Fig. 6A. We found that $\Delta d$ increases with the distance between the city centers. For instance, in the case of a five-center city such that the four new centers are established at a distance of 12km outward from the original center, $\Delta d$ reaches $22.3\%$, suggesting that new urban centers should be established in peripheral areas to achieve a more effective reduction in commuting distance through home swapping. This suggests a strategic approach to urban planning whereas developing multiple activity centers with a distributed pattern and a larger distance between them can significantly improve commuting efficiency by home swapping, hence achieving a more efficient and energy-saving city. Similar result can be obtained in Shenzhen and Abidjan as shown in Fig. S19B and Fig. S23B respectively.
We further select bustling areas in the four peripheral counties of Shijiazhuang as new activity centers, to explore the impact of establishing multiple peripheral centers on the reduction in commuting distance. The results in Fig. 6B show that as the number of activity centers increases, the reduction of commuting distance increases; from a single center to dual centers, even without the corresponding supporting amenities, there is a significant reduction in commuting distance after SDGHS. From dual centers to multiple centers, the marginal increase in the reduction of commuting distance slows down. Other than the reduction in the average commuting distance, we show in Fig. 6C the congestion coefficient decreases with the number of activity centers, while carbon emissions are reduced after SDGHS under medium tolerance. This indicates that home swapping with the establishment of a polycentric city is a promising urban-planning solution in alleviating congestion and reducing carbon emissions.
Finally, we further considered enlarging the amenity accessibility of the newly added activity centers and increasing the housing vacancy based on the original city center (see Fig. S12); conclusions similar to the above were obtained.
III Discussion
In this paper, we explored the benefit brought by an innovative approach to swap homes between households to reduce traffic congestion and carbon emissions in urban areas. By analyzing over 400,000 mobility trajectories in the Chinese city Shijiazhuang, we found that the overall city-wide commuting distance can be potentially reduced by over $50\%$. This finding is significant as it underscores the interplay between the distributions of citizens’ residence, workplace locations and urban traffic congestion. In addition, due to both a shorter commuting distance and reduced urban congestion, carbon emissions are also greatly reduced, which can be another important impact of home swapping. Our study is the first to verify the benefit brought by home-swapping using such a big real data.
Interestingly, to make home swapping more realistic, our analysis also revealed that when households considered the distance to the city center, housing price and amenity accessibility in swapping home, the reduction in commuting distance declines but still accounts to $13\%$, which is a substantial decrease. Moreover, among these factors, the distance from the city center emerged as the most influential factor for urban residents in choosing their residence location and whether to swap home. These results support the coordinated allocation of residence as an innovative strategy to substantially alleviate peak-hour traffic congestion and carbon emissions in densely populated metropolitans.
Though complete mandatory home swapping may not be readily feasible in reality, our findings are crucial in the broader context of urban planning and sustainable city development. By demonstrating the strong impact of residence allocation on traffic status, our work provides valuable insights and guidance for urban planners and policymakers in urban planning. It stresses the need for a holistic consideration on urban design, where residential planning is not only about housing city residents but also about integrating socio-economic factors including amenity accessibility, financial and environmental considerations. This study contributes significantly to the existing body of knowledge by providing empirical evidence on a unique, innovative yet determined home swapping approach on how urban structure can be coordinated and reformed, specifically towards a polycentric model to save energy, resources and time by significantly minimizing the fundamental source of traffic congestion. This is particularly relevant to tackle the challenges of both rapid urbanization and increasing urban population, in a board sense sheds light on sustainable city development.
While this study offers important and interesting insights, it has limitations. First, although the major analysis was based on a large dataset obtained in Shijiazhuang and similar results have been observed by using data in Shenzhen and Abidjan, analyses on more cities with different cultural, economic, or geographical characteristics can further extend its generalizability. Furthermore, the study primarily assumed residents commute by motor vehicles, not taking into account other modes of public transportation such as buses or metro. Future research could extend this work by examining a more diverse range of cities, incorporating multi-modal transportation data, and exploring the interplay between residence locations and different types of commuting behaviors. Additionally, investigating the psychological and social impacts of home swapping on individuals and communities would provide more understanding on this innovative approach, its feasibility and implications on urban planning and city development.
Materials and Methods
Dataset of Mobility Tracks. Our original data is composed of 9-day 4G communication records between base stations and mobile phones served by one of the three major service providers from May $22^{\rm nd}$ to $28^{\rm th}$, 2017, and from June $3^{\rm rd}$ to $4^{\rm th}$. The original data contain more than 3 million users out of a total population of 10 million in Shijiazhuang, a major city in northern China. There are over 11500 base stations throughout the city, and the position of an individual is recorded as the location of the nearest base station as long as his/her mobile phones communicate with the base stations in 4G. As some phone applications constantly exchange data with back-end servers, the position of individuals can be recorded up to a high frequency of every second. We then divide each single day into $96$ time windows, each with a duration of 15 minutes, and consider a user stops in a location if he/she stays there continuously or discontinuously for 10 minutes or more within the 15-minute time window. Otherwise, if the user does not stay in the same location for more than 10 minutes within a time window, we identify the data point as “moving”.
Identifying Individual Residences and Workplaces in Shijiazhuang. A user’s home location is defined as the place where they spend the longest time between 8 PM and 7 AM, and the total duration of stay each night must exceed 6 hours. If these criteria are not met, the location of the home for that individual on that day is considered unknown. The workplace is defined as the location outside the individual’s home, where a he/she spends the most time between 7 AM and 8 PM, and the total duration of stay each day must exceed 2 hours. If these requirements are not met, the workplace data for that individual on that day is empty. Given five weekdays of data, a maximum of five home location assessments can be made for one individual. To mitigate the impact of base station spacing, base stations with home locations less than 250 meters apart are merged, using the location where the longest stay occurred as a substitute. Then, the individual’s nighttime data is updated to obtain the position of his/her home each night. For an individual with five days of home location assessments, the location is considered a clear residential characteristic only if at least four days point to the same place. Similarly, there are five records for the workplace. We define the workplace for an individual as the location that appears most frequently among the five assessments. If two locations occur with the same number of times, the total duration of stays over the five days at both locations is compared, with the longer duration determining the individual’s workplace.
In our study, considering that urban congestion is primarily concentrated in the central areas of the city, we focus on the main urban districts in Shijiazhuang, including Yuhua, Xinhua, Chang’an, Qiaoxi, Zhengding, Gaocheng, Luquan and Luancheng. Only individuals live or work within these main urban districts are considered, resulting in a trajectory dataset from 422,454 anonymous commuters.
Similarity between Individual Trajectories. The daily similarity $S_{ij}$ between individual $i$ and $j$ is given as follows,
$S_{ij}=\frac{1}{L_{ij}}\sum_{t=1}^{M}\Theta(\theta_{c}-\theta_{ij}(t))\Theta(d_{c}-d_{ij}(t))$ (2)
where the step function $\Theta(x)=1$ if $x\geq 0$ or otherwise $\Theta(x)=0$. The variable $\theta_{ij}(t)$ is the angle between the residence and the position of individual $i$ and $j$ at time $t$; $\theta_{c}=30^{o}$, while $d_{ij}$ represents the distance between individuals $i$ and $j$ at time $t$ and $d_{c}=272m$, which is the average street length in Shijiazhuang obtained from OpenStreetMap (see SI section 1). $M=96$ is the total number of time intervals in one day, and $L_{ij}$ represents the number of time intervals when the location information of both individuals $i$ and $j$ are available and they were not at their home. We require $L_{ij}\geq 4$ for $S_{ij}$ to be considered.
The KL Divergence between the Spatial Distributions of Residence and Workplace. The KL Divergence between spatial distributions of residence and workplace is calculated as follows,
$KL(p||q)=\sum_{x}p_{x}\log\frac{p_{x}}{q_{x}}$ (3)
where $p_{x}$ and $q_{x}$ denote the density of residence and workplace in grid $x$, respectively. A smaller KL divergence indicates a smaller degree of difference between the two distributions.
Inferring Commuting Route and Home-Departure Time. To obtain the actual commuting distance before and after commuters swap their homes, we combined different residence and workplace locations in pairs and crawled the routes for each Origin-Destination (O-D) pair via the navigation function on Baidu Map. Due to the large volume of data, we grouped different locations to the nearest road center. If the residence and workplace are on the same road, we used the same location of origin. Each data entry includes the residence, workplace, commuting distance, commuting route, and the time spent on each section of the route.
The inferred home-departure time is the average time each commuter first leaves home between 5 AM and 11 AM on the 5 weekdays covered by the data. Commuters not meeting this criterion are assigned departure times proportionally. As shown in Fig. S2 of the SI, the proportion of commuters departing home between 7 AM and 8 AM is the highest. The home-departure time after home swapping is inferred as the time the commuter has to start their journey in order to reach their workplace at the same time before home swapping. The distributions of home-departure time after GHS and SDGHS are shown in Fig. S3A and S3B of the SI.
Average Congestion Coefficient. To explore the impact of shortened commuting distances on traffic congestion after home-swapping, we define an average congestion coefficient $I_{t}$ as follows:
$I_{t}=\bar{c_{\alpha}^{t}}=\frac{\sum_{\alpha=1}^{M_{t}}c_{\alpha}^{t}}{M_{t}},$ (4)
where $M_{t}$ is the number of roads in the set of congested and slow-moving segments at time $t$ in empirical situation, and $c_{\alpha}^{t}$ is defined as the congestion coefficient for each road segment de2014determining , given as follows
$c_{\alpha}^{t}=\frac{N_{\alpha}^{t}}{\tilde{N}_{\alpha}},$ (5)
where $N_{\alpha}^{t}$ is the number of vehicles traveling in the same direction on road segment $\alpha$ at time $t$, and we calculate the start time of commuters on each segment based on their travel times and departure times. $\tilde{N}_{\alpha}$ is the maximum traffic volume of the road $\alpha$, given by
$\tilde{N}_{\alpha}=\frac{1000\times\tilde{v}_{\alpha}}{l_{c}+l_{a}^{\alpha}},$ (6)
which varies depending on the speed limit $\tilde{v}_{\alpha}$ of road segment $\alpha$(see navigation data in SI Section 1). $l_{c}$ is the average length of a vehicle, i.e. 4 meters; $l_{a}^{\alpha}$ is the length of the safety distance between vehicles, which is dependent on the maximum driving speed of road $\alpha$, e.g. $l_{a}^{\alpha}=100m$ if speed limit $\tilde{v}_{\alpha}=100km/h$ mcshane1990traffic .
The higher the average congestion coefficient, the more congested the city is. It is also important to note that the coefficient is calculated hourly from data in four 15-minute intervals, and commuters using non-motorized vehicles have been excluded. If a commuter passes road $\alpha$ during time interval $t$, then $N_{\alpha}$ at time $t$ increased by one.
Identifying Housing Price. We identify the price of a housing unit at a residence location based on their price listed in the second-hand market from the website Lianjia.com for Shijiazhuang, where information such as the name of the properties, the latitude and longitude of the residential communities, their total price and the unit price, are available (see SI Section S1). The city was divided into a grid of $500\times 500$ squares, and residences in the second-hand market, i.e. houses, apartments, etc., were placed in the corresponding grid based on their latitude and longitude. The average housing price of a residence in each grid is calculated by averaging the prices of all the second-hand residences within the grid. For grids without housing price data, the price of a housing unit is assumed to be the average of the housing prices of the nearest four grids.
Inferring the Dissimilarity in Amenity Accessibility. We utilize the data of point of interest (POI) crawled from Gaode, a map app used in China, which includes each POI’s name, type, location, area, and street information. We then select seven major categories of amenities: (1) dining, (2) shopping, (3) transportation facility, (4) science, education and culture, (5) services, (6) sport and leisure, and (7) medical and security, such that the weighted proportion of the $k$-th type of amenities at location $x$ is given as follows,
$a_{x}^{(k)}=\frac{n_{x}^{(k)}}{\sum_{k=1}^{7}n_{x}^{(k)}}\times\frac{N^{(k)}}{\sum_{k=1}^{7}N^{(k)}},$ (7)
where $n_{x}^{(k)}$ and $N^{(k)}$ denote the number of the $k$-th type of amenities, at base station $x$ and across all studied locations, respectively. On top of the amenities right at their residential location, residents can utilize amenities nearby. We thus define the amenity accessibility of the $k$-th type of amenities at location $x$ as follows,
$A_{x}^{(k)}=\sum_{y}a_{x}^{(k)}\Theta(d_{xy}-L)\left[1-\left(\frac{d_{xy}}{L}\right)\right]^{2},$ (8)
where $\Theta(x)$ is again the step function such that $\Theta(x)=1$ when $x\leq 0$ or otherwise $\Theta(x)=0$; $L$ is the average street distance in Shijiazhuang, which is determined from Open Street Map (OSM) data to be $L=272m$. Finally, we calculate the residential similarity between location $x$ and $y$ by using the cosine similarity between $\vec{A}_{x}$ and $\vec{A}_{y}$, where $\vec{A}_{x}$ is the vector composed of elements $A_{x}^{(k)}$. The amenity dissimilarity $D_{xy}$, which describes the difference of the accessibility to various amenities between location $x$ and $y$, is given by
$D_{xy}=1-\frac{\vec{A}_{x}\cdot\vec{A}_{y}}{|\vec{A}_{x}||\vec{A}_{y}|}$ (9)
Urban Carbon Dioxide Emissions For exploring how GHS affects carbon dioxide emissions from the whole city, we define the carbon dioxide emission $E(t)$ during each time interval $t$ as in kissinger2019detailed :
$E(t)=\sum_{i=1}^{N_{c}}\sum_{\alpha\in Path_{i}(t)}d_{i}^{\alpha}\times\zeta\times c_{\alpha}\times S_{\alpha}\times R_{\alpha},$ (10)
where $N_{c}$ is the number of commuters traveling at time interval $t$; $Path_{i}(t)$ is the set of road segments commuter $i$ passes during time interval $t$, and $d_{i}^{\alpha}$ is the length of road segment $\alpha$; $\zeta$ is the carbon dioxide emission factor, which is equal to 0.209 $kgCO_{2}e/km$; $c_{\alpha}$ is the number of vehicles passing road $\alpha$ at the same time with commuter $i$; $S_{\alpha}$ and $R_{\alpha}$ represent the slope and the type of road $\alpha$, which are both set as constants for simplicity. In addition to the total carbon emissions, we can also obtain the emissions from each commuter or on each road in each time interval.
Acknowledgments.
C.Z. acknowledges the Natural Science Foundation of Hebei (No. F2020205012), the Youth Top Talent Project of Hebei Education Department (No. BJ2020035), and the Project Supported by Science Foundation of Hebei Normal University (No. L2023K04). The work by C.H.Y. is supported by the Research Grants Council of the Hong Kong Special Administrative Region, China (Projects No. EdUHK GRF 18301217, and No. GRF 18301119), the Dean’s Research Fund of the Faculty of Liberal Arts and Social Sciences (Projects No. FLASS/DRF 04418, No. FLASS/ROP 04396, and No. FLASS/DRF 04624), and the Internal Research Grant (Project No. RG67 2018-2019R R4015 and No. RG31 2020-2021R R4152), The Education University of Hong Kong, Hong Kong Special Administrative Region, China.
Author contributions.
C.Z., C.H.Y. and A.Z. designed the research, C.Z., Y.L., X.H. and J.D. performed the experiments, C. Z., C.H.Y. and A.Z. analyzed the data, C.H.Y. and A.Z. wrote the paper.
Competing financial interests. The authors declare no competing financial interests.
Data and materials availability. The raw data cannot be shared by the authors due to the Mobile Privacy Policy of China, but might be available upon reasonable request submitted to the mobile network operator (China Mobile) at http://it.10086.cn/services/##\##dsjArea.
Supplementary Information
Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions
Chen Zhao, Yuqing Liu, Xiaoyue Hou, Jianghui Ding, Chi Ho Yeung, and An Zeng
IV Data Description
The details of the 12 independent datasets used in this study are given as follows:
  • Shijiazhuang Individual Mobility Trajectory Dataset [1]: The empirical mobility trajectory big data used in this article is based on 4G communication records between cell towers and mobile phones provided by one of the three major service providers in Shijiazhuang, a major city in northern China, from May 22 to 28 and June 3 to 4, 2017. This dataset covers a total of 9 days (5 weekdays and 4 weekend days). To identify meaningful locations in users’ mobility trajectories, the dataset was divided into 15-minute intervals, 96 intervals per day. If an individual stays at the same location for more than 10 minutes in one 15-minute interval, the user is considered to have stayed in that location.
  • Shenzhen Living-Working Dataset [2]: The Shenzhen data is a dataset collected by Zhilu Yuan and others, representing the largest urban living-working tracking dataset from the Shenzhen 2015 Census logs. It includes records of living and working locations for 5 million anonymous commuters, covering 71 geographically adjacent subdivisions across 10 districts. The locations of each commuters’ residence and workplace are indicated by the latitude and longitude of their respective streets.
  • Abidjan Call-Detail Record (CDR) Dataset [3]: The Cote d’Ivoire CDR data provides anonymous user IDs, timestamps, and sub-county IDs (including their locations). When a user makes a call from a new location, the user’s movement can be identified.
  • Shijiazhuang Commuting Route Navigation Data: To obtain the actual commuting distances before and after commuters swap their homes, first, different residential and workplace locations are paired. Due to the large volume of data, in this study we aggregate different location points to the nearest road center. If the residence and workplace are on the same road, the original location points are used. In this study, we used the aggregated locations to crawl the driving routes for each pair of commuting locations on the Baidu map [4], during morning traffic peak (6:00 A.M.-10:00 A.M.) and early evening traffic peak (4:00 P.M.-8:00 P.M.) in weekdays, which includes the time taken for an individual to travel through each road segment, the type and the distance of the road segment, and navigation details. The final data include the individual’s starting location, the aggregated starting location, commuting distance, commuting time, and commuting route.
  • Shenzhen Commuting Route Navigation Data: The residential-workplace dataset of Shenzhen is a street-level record, and we represent the whole street using the road center’s latitude and longitude, and crawl the driving routes for each pair of the commuting locations on the Baidu map.
  • Shijiazhuang On-demand Vehicle Dataset: The data from DiDi on-demand vehicle in Shijiazhuang for five weekdays in July 2017, including the pick-up and drop-off location of each trip.
  • Shijiazhuang Housing Price Data: The information of second-hand housing prices from Lianjia.com [5] within the Shijiazhuang city area in 2017, including total price, unit price, community name, and community location.
  • Shenzhen Housing Price Data: The information of second-hand housing prices from Lianjia.com [6] and Anjuke.com [7] within the Shenzhen city area in 2015, including total price, unit price, community name, and community location.
  • Shijiazhuang Point-of-interest (POI) Data: The information on various types of Points of Interest (POI) within the Shijiazhuang city area collected from Gaode in 2017, including store name, type, location, area, and street information.
  • Shenzhen POI Data: Information on various types of Points of Interest (POI) within the Shenzhen city area collected from Gaode in 2015, including store name, type, location, area, and street information.
  • Shijiazhuang 2017 Census Data and Shenzhen 2015 Census Data [8].
  • Shijiazhuang Urban Road Network Data: Obtained from OpenStreetMap [9] to facilitate aggregating residence or workplace locations to the nearest road center. The total number of road segment is 28.9 million. The average street length in the city, calculated from crawling the length of streets at all levels, is found to be 272 meters.
V Statistical Analysis Based on Data
Identifying Commuters from the Same Household. To obtain the actual number of commuter households from the data, we take reference of the 7-th Census of Shijiazhuang, revealing the proportion of households from one-person to six-person to be 0.12, 0.27, 0.23, 0.19, 0.097, and 0.093 respectively. According to the demographic yearbook, one commuter can support up to 1.92 people, hence 422,454 commuters can support a maximum of $N_{p}=811112$ people. Denoting $p_{i}$ is the proportion of households with $i$ members, the number of households with $i$ members is $n_{i}=p_{i}\times N$, where $N$ is the total number of households. The number one-person to six-person households are found to be 30,870, 69,458, 59,168, 48,878, 24,953, and 23,924 respectively.
Next, we assumed that the number of commuters in one household is just the supply minimum $n_{c}^{min}$, or $n_{c}^{min}-1$, without considering supply redundancy. Therefore, the potential number of commuters in households with different number of members is shown in the following table:
family size 1 person 2 persons 3 persons 4 persons 5 persons 6 persons
0 commuter $\surd$
1 commuter $\surd$ $\surd$ $\surd$
2 commuters $\surd$ $\surd$ $\surd$ $\surd$
3 commuters $\surd$ $\surd$ $\surd$
4 commuters $\surd$
Table 1: The possible commuter families in different household types.
Thus, the number of households with different number of commuters can be calculated, taking two-person households as examples:
$\frac{N_{p}^{2}}{N_{c}^{2}}=\frac{2n_{2}}{n_{2}^{1}+2n_{2}^{2}}=1.92$ (11)
$n_{2}^{1}+n_{2}^{2}=n_{2}$ (12)
where $N_{p}^{2}$ is the total number of people in $2$-person households; $N_{c}^{2}$ is the total number of commuters in $2$-person households ; $n_{2}^{j}$ is the number of $j$-commuter families in $2$-person households. Thus, the actual number of different commuter households in the data can be calculated, with the number of households from single-commuter to four-commuter households to be 108,515, 90,843, 40,076, and 3,006 respectively.
Commuter Households Obtained through Trajectories Similarity. Since individual mobility trajectory data do not include the information about household membership, we used data from the 4 days in the 2 weekends to identify commuters from the same household. According to the census information, most of households in Shijiazhuang have less than $7$ members, and each working person supports up to 1.92 individuals. Therefore, we assume that each household has a maximum of 6 members and 4 commuters. We identify family members by calculating the similarity between the mobility trajectories among individuals who share the same residence location. By analyzing the distribution of similarities $S_{ij}$ between individuals, those with $S_{ij}>0.5$ are considered household members. An individual might have similarities over 0.5 with multiple people; these similarities are sorted from the highest to the lowest, and only a maximum of five individuals with the highest similarity are considered to be potential household members. Ultimately, the ratio of single-, double-, triple- and quadruple-commuter households was found to be 65:26:6:2.
Compared to our census-deduced household ratios of 37:31:14:1, double and triple-commuter household are under-estimated, while single and quadruple-commuter households are over-estimated. To match the constructed and the deduced household ratio, we randomly selected 3,302 quadruple-commuter households, using the three members with the highest similarity in each household to form triple-commuter household, with the remaining person form a single-commuter household, resulting in 192,777 single-, 76,533 double-, 21,529 triple-, and 3,006 quadruple-commuter households. In this case, the number of single-commuter households still exceeded the inferred amounts, due to the criteria of $S_{ij}\geq 0.5$, which is a more stringent criteria for identifying house members, and thus leads to an overestimation of single-commuter households.
To identify households with a ratio in agreement with the census of Shijiazhuang, we release the stringent criteria and form households among individuals from single-commuter households with $S_{ij}\geq 0.1$. We finally obtained a ratio of households matching the deduced ratio, such that single to quadruple-commuter household to 108,515, 90,843, 40,076, and 3,006 respectively.
Using the shortest travel distance by motor vehicle. The mode of transport is highly related to the commuting distance, we analyzed the distance for on-demand motor vehicles the Didi in China in 2017 (as shown in Figure S1)to identify a threshold for urban residents to differentiate between short and long distances. As we can see in Fig. S1, the number of on-demand travel peaks at a distance of 2.5 km. Therefore, we assume that 2.5 km is the shortest distance that urban residents opt for motor vehicle travel in their daily commuting travel between their home and workplace.
Supplementary Figure S1. The Distribution of Travel Distance of DiDi On-demand Vehicles in China.
Empirical Commuter Departure Times.
Supplementary Figure S2. The Distribution of Empirical Time of Departure from Home and Workplace. Panel A shows the empirical time of departure of commuters from their home during the morning rush hour; Panel B shows the empirical time of departure from their workplace during the evening rush hour.
Commuters’ empirical departure time is set as the average time of their department from their home between 5 AM and 11 AM on the 5 weekdays in the dataset. The return time from the workplace for commuters is set as the average time of their departure from their workplace between 4 PM and 11 PM on the 5 weekdays in the dataset, as shown in Figure S2. Commuters departures are notably observed during the morning rush hour between 7-8 AM (Fig. S2A), and most commuters returns are observed during the evening rush hour between 5-6 PM (Fig. S2B), though the return time distribution is noticeably more dispersed.
VI Greedy Home Swapping
Change in Departure Times. The empirical distribution of commuters’ departure time is shown in Figure S2A. As we can see, commuters departure times peak between 7-8 AM, with the highest number of departures at 7:30 AM. It is stipulated that if the morning arrival time to the workplace and the evening departure time from the workplace remain unchanged, the new morning departure time from their new home and the evening new arrival time back to their new home after GHS would shift and can be deduced from navigation data.
The distribution of the commuters’ new departure time from their new home after GHS is shown in Figure S3A. The results revealed that the departure time becomes more dispersed compared to that in the empirical data, with fewer residents departing during peak time (7:30 AM). The distribution of the new departure times after SDGHS with medium tolerance is shown in Figure S3B, which is similar to that after GHS, but the number of commuters who use motor vehicles increases, resulting in a lower reduction in commuting distance, i.e. $\Delta d$.
Supplementary Figure S3. The Distribution of Home-Departure Times During the Morning Rush Hour. Panel (A) shows the departure time after GHS; Panel (B) shows the departure time after SDGHS under medium tolerance.
Snapshots of congestion status obtained based on the empirical data and after RHS and GHS during rush hour. By computing the congestion coefficient of each road $c_{\alpha}$ (see Materials and Methods in main text) during the morning peak hours between 7-8 AM and between 8-9 AM , and the evening peak hour between 5-6 PM obtained by the empirical data, and after RHS and GHS, we visualize the congestion coefficient in Figure S4. We found that the road congestion after RHS is greater than that based on the empirical data at all time, while the road congestion after GHS is less than that based on the empirical data. Moreover, after RHS , most road segments are congested, whereas after GHS, only a few segments remain congested, with most becoming slow-moving.
Supplementary Figure S4. The congestion coefficient obtained based on the empirical data and after RHS and GHS. The top three maps show the congestion coefficient during the morning peak hours between 7-8 AM; the middle three maps show the congestion coefficient during the morning peak hours between 8-9 AM; the bottom three maps show the congestion coefficient during the evening peak hours between 5-6 PM.
Average Congestion Coefficient of City Affected by Home Swapping.
Supplementary Figure S5. The Average Congestion Coefficient of Shijiazhuang over Rush Hour.
Empirical data show that the average congestion coefficient is highest during the morning rush hour (7-8 AM) and the evening rush hour (5-6 PM). The average congestion coefficient after RHS is significantly higher than that based on the empirical data; whereas $I_{t}$ after GHS is much lower than that based on the empirical data (see Figure 3C in the main text). $I_{t}$ is similar under medium tolerance on amenity accessibility and housing price, but after SDGHS under the medium tolerance, the city is more congested and closer to empirical data, as shown in Figure S5.
Supplementary Figure S6. The number of congested, slow-moving, and free-flowing road segments during the morning rush hour between 7-8 AM.
Since the average congestion coefficients $I_{t}$ after GHS under different factors do not vary significantly, we further analyze the changes in the number of congested roads with $c_{\alpha}\geq 1$, slow-moving roads with $10.4\leq c_{\alpha}<$ 1, and free-flowing roads with $c_{\alpha}<0.4$ during the morning rush hour between 7-8 AM under different restrictive factors of GHS, as shown in Figure S6. After GHS, there is a noticeable decrease in congested and slow-moving road segments, while free-flowing road segments significantly increase. By considering only the amenity accessibility (a-GHS), housing price (p-GHS) and the distance to the city center (c-GHS), as well as all the three socio-demographic factors (SDGHS), the average congestion coefficient $I_{t}$ gradually increases, the number of severely congested and slow-moving road segments gradually increases, while the number of free-flowing road segments decreases. Although the number of congested roads in a-GHS is lower than that after GHS, slow-moving segments increase, and free-flowing road segments decrease. Additionally, $I_{t}$ after p-GHS is similar to that after a-GHS, but with a higher number of congested and slow-moving segments.
Considering dual socio-demographic factors (medium tolerance) during home-swapping. We also consider $\Delta d$ by considering two out of the three socio-demographic factors in GHS, as shown in Figure S7. In this discussion, we use the medium tolerance SDGHS discussed in Figure 4 of the main text. The reduction in commuting distance, i.e. $\Delta d$ under (1) the joint consideration of the distance to the city center and amenity accessibility (c&a-GHS) is 23.6%, (2) the joint consideration of of the distance to the city center and housing price (c&p-GHS) is 26.6%, and (3) the joint consideration of housing price and amenity accessibility (p&a-GHS) reaches 36.29%. The results show that the distance from the city center has a significant impact on the benefit brought by GHS, followed by amenity accessibility, and finally the housing price.
Supplementary Figure S7. The Reduction of the Average Commuting Distance, i.e. $\Delta d$, under the joint consideration of two out of the three socio-demographic consideration.
The Impact of the Willingness to Swap Home. In GHS, we assume that all households are willing to swap their home as long as the total commuting distance of the two swapping households decreases, even at the expense of an increase in the commuting distance of one of them. Here, we assume that the households with an increase of commuting distance has an $x\%$ willingness to swap but $1-x\%$ to reject.
Supplementary Figure S8. $\Delta d$ after GHS or after SDGHS (medium tolerance) considering households’ willingness to change residences.
As different households have different levels of willingness to swap homes, $\Delta d$ after GHS method with different willingness $x\%$ is shown by the blue line in Figure S8. The results show that $\Delta d=40.9\%$ when households do not swap residences if the average commuting distance of one household increases. As the willingness to swap increases from 0% to 25%, there is a noticeable increase in $\Delta d$. However, from a willingness of 25% onwards, the increase in $\Delta d$ slows down as the willingness of swapping $x\%$ increases.
Moreover, we consider households willingness to swap home in SDGHS under medium tolerance and analyzes $\Delta d$ after home swapping under different willingness $x\%$, as shown by the purple line in Figure S8. The results find that as the $x\%$increases, $\Delta d$ also increases. When $x=0$, $\Delta d$ is 9.6%; when $x>=25$, there is no significant increase in $\Delta d$.
VII The Reduction of Carbon Emissions.
Supplementary Figure S9. The Road carbon emissions obtained based on the empirical data and after RHS and GHS. The top three diagrams show the CO2 emissions on the roads during the morning peak hours between 7-8 AM; the middle three diagrams show the CO2 emissions during the morning peak hours between 8-9 AM; the bottom three diagrams show the carbon emissions during the evening peak hours between 5-6 PM.
We computed and visualized the carbon dioxide emissions of each road segment during the morning peak hours between 7-8 AM, between 8-9 AM, and the evening peak hour between 5-6 PM based on the empirical data and after RHS and GHS , as shown in Figure S9. We found that the carbon emissions under RHS at are greater than that based on the empirical data at all time, while those after GHS are significantly less than that of the empirical data. The carbon dioxide emissions between 7-8 AM during the morning peak hours are higher than those between 5-6 PM during the evening peak hours. Additionally, after RHS, most road segments emit a large amount of CO2, whereas after GHS, only a few segments have higher emissions, and most road segments emit a small amount of CO2 (road segments emitting 0.4 tons of carbon are marked in red, and those emitting between 0.2 and 0.4 tons are in yellow, and those emitting less than 0.2 tons are in green).
Supplementary Figure S10. The carbon emissions during rush hour under different conditions.
The reduction in carbon emissions after GHS, and SDGHS is shown in Figure S10 for each hour during the peak commuting hours. In the morning peak hours, the average commuting distance after home swapping become shorter for most households. After home swapping, each individual is supposed to arrive at their workplace at the same time before home swapping, which leads to a dispersed departure time (see Figure S3). It was found that the reduction in carbon emissions decreases over time. Similarly for the evening peak hours, where each individual is supposed to leave workplace at the same time after home swapping, the shorter commuting distances lead to a gradual increase in carbon emissions reduction over time. Among all the cases, the GHS emissions reduction is the highest, and the emissions reduction of SDGHS is the lowest.
Supplementary Figure S11. The heatmap of changes in individual emissions before and after GHS.
We further analyze the changes in commuters’ emissions before and after residential allocation by GHS, as shown in Figure S11. The individual’s unaltered emissions following GHS is indicated by the black dashed line, and the majority of commuters experience a reduction in emissions after changing their residence, as changes in residence lead to non-motorized travel, reducing emissions to zero. However, there are cases where individuals with initially short commuting distances experience an increase in both commuting distance and emissions after changing residence.
VIII Adding new activity centers based on the original city center.
Supplementary Figure S12. $\Delta d$ after adding new activity centers, where the green line and symbols represent the addition of the corresponding amenity configurations, and the blue line and symbols represent the addition of the corresponding vacant housing.
(1) Adding new activity centers and the corresponding amenity configurations. In this study, different numbers of new activity centers are established in surrounding county towns, and the amenity configurations around the new centers are changed accordingly. Based on the distance of each base station from the original city center and the proportions of various amenities, as represented by POI, around each base station, the corresponding POIs are added to the surrounding of the new centers, as shown in Figure S12. The yellow line corresponds to an increase of the number of centers without the corresponding amenities, while the green line represents adding the corresponding POI configurations around the new activity centers. Comparing the two lines, there is not a significant increase in $\Delta d$ with the corresponding amenity configurations added.
(2) Adding new centers and increasing vacant housing. Without providing new workplaces, the residential capacity of the new centers is proportionally increased as in the original city center. As shown by the blue line in Figure S12, as the number of centers increases, $\Delta d$ after applying medium tolerance SDGHS gradually increases, showing an improvement compared to just adding new centers.
IX Impact of Commuting Mode Shift Threshold.
After GHS, there are changes in the spatial configuration of residences and workplaces as well as commuters’ commuting distances, which also lead to the corresponding changes in commuters’ modes of transportation. Through analyzing the travel distance of on-demand Didi vehicles in China, we assume 2.5 km to be the shortest distance for commuters to travel by motor vehicle. It is considered that a commuter who travels less than 2.5 km does not contribute to traffic congestion and carbon emissions, and instead uses non-motorized transportation. First, we studied the impact of GHS under different restrictive conditions on the changes in commuters’ modes of transportation, as shown in Figure S13. The results show that after residence re-allocation by GHS, 64% of commuters used non-motorized transportation, which is significantly higher than the empirical number of non-motorized commuters (36.5%). Under medium tolerance with three combined factors, 44.1% of commuters used non-motorized transportation, which is only a 7.6% increase compared to the empirical data.
Supplementary Figure S13. The proportion of non-motorized vehicles used by commuters, after different versions of GHS under medium tolerance.
While we have assumed 2.5 km to be the shortest distance for commuters to travel by motor vehicles, we also examined other thresholds of assumed minimum distance at $2$, $1.5$, $1$, $0.5$, and 00 km, as shown in Figure S14. As the threshold decreases, the proportion of non-motorized vehicles shows a significant decline. Under medium tolerance with SDGHS, 9% of commuters live and work within a very short commuting distance.
Supplementary Figure S14. The proportion of non-motorized vehicles used by commuters, under different assumption thresholds of minimum distance in the use of motorized vehicle, after different version of GHS.
In addition, we also explore the impact of the assumed thresholds of the minimum distance of the use of non-motorized vehicles on the average congestion coefficient and carbon emissions. Figure S15A shows the impact of different values of thresholds on traffic congestion during the 7-8 AM peak hours. As the threshold value decrease, the number of used motor vehicles increases, the occupancy of road sections during peak hours increases, thus increasing congestion. When all commuters travel by motor vehicles, congestion also increases. By considering the three socio-demographic factors, there are fewer non-motorized vehicles, and the change in the minimum travel distance for using motor vehicles does not significantly impact traffic congestion.
On the other hand, for carbon emissions, as the threshold decreases, the number of motor vehicles increases, and the carbon emissions increase, but not significantly, as shown in Figure S15B. It is because the reduction in the minimum travel distance for using motor vehicles increases the usage of motor vehicles commuting over a short distance, which does not significantly impact carbon emissions.
Supplementary Figure S15. The impact of the threshold values of the usage of the motorized vehicles. (A) shows the dependence of the average congestion coefficient $I_{t}$ on the threshold values. (B) shows the dependence of the reduction of emission reductions on the threshold values.
X Examples of home swapping in cities other than Shijiazhuang
Home Swapping in Shenzhen. Similar to Shijiazhuang, we first analyzed the spatial distributions of residence and workplace of commuters in Shenzhen’s urban area. Figure S16 shows the heat map of residence and workplace locations in Shenzhen. Among them, Futian, Bao’an, and Luohu Districts are the most bustling areas, with the highest number of residences and workplaces represented by the red areas, while Longgang , Yantian, and Pingshan Districts represented by the blue and green areas on the right have fewer residences and workplaces. By the computing the KL divergence between the two spatial distributions, we found a KL divergence value of 0.114, indicating that the distribution of residences and workplaces is very similar across the city, suggesting a large potential benefit brought by GHS.
Supplementary Figure S16. The spatial distributions of residence and workplace in the urban area in Shenzhen. (A) shows the heat map of residences in Shenzhen; (B) shows the heat map of workplaces in Shenzhen.
Even though the spatial distributions of residence and workplace in Shenzhen is very similar, the city still experiences traffic congestion [10], indicating that long commuting distance of commuters may be a reason. Therefore, we analyzed and showed the distribution of commuting distance of Shenzhen’s commuters in Figure S17. As we can see, the peak commuting distance for most commuters exceeds 10 km, with more than half of the commuters experiencing long-distance commutes daily, suggesting a significant room for exploitation. It is worth noting that the commuting distances used here are the distances obtained by navigation devices.
Supplementary Figure S17. The distribution of commuting distance in Shenzhen.
Due to the limitation of the Shenzhen’s data without detailed individual trajectory information, it is impossible to infer household membership through individual trajectory analysis. Therefore, we constructed single and double households with a ratio of 6:4 based on statistics from the Shenzhen 2015 City Population Census Yearbook. Initially, commuters living at the same latitude and longitude in the data were randomly paired to form double-commuter families, from which 2806394 individuals were randomly selected to form double-commuter households, leaving the remaining commuters as single-commuter households. Results show that $\Delta d$ after GHS reached 62%.
Supplementary Figure S18. $\Delta d$ obtained after considering single socio-demographic factor in Greedy Home Swapping (GHS) in Shenzhen.
Next, we used Shenzhen’s data on amenity (i.e. point of interest (POI)), housing price, and commuter routes based on the output from navigation devices to analyze the socio-demographic GHS, as shown in Figure S18. From Figures S18A and B, it can be observed that after c-GHS and p-GHS, $\Delta d$ improves as the restrictions are relaxed. However, changes in the dissimilarity in amenity accessibility in a-GHS did not significantly enhance $\Delta d$ in Figure S18C, likely due to data granularity, with data spatial resolution at the road level and POIs calculated around road surroundings, leading to indistinguishable road in some cases. In addition, as we see in Figure S18B, there is a notable increase in $\Delta d$ when the tolerance on the difference of housing price increases from 10% to 30%, which aligns with the wide range of housing prices in Shenzhen.
We also examine the joint consideration of all the socio-demographic factors in GHS in Shenzhen similar to the case of Shijiazhuang in the main text. As shown in Figure S19A, under low tolerance ($T_{center}=1km$, $T_{price}=10\%$, $T_{amenity}=0.1$), $\Delta d$ is less than 10%. As tolerance increase, $\Delta d$ noticeably improves, consistent with the conclusions in Figure 4D of the main text. However, under high tolerance ($T_{center}=5km$, $T_{price}=50\%$, $T_{amenity}=0.5$), $\Delta d$ in Shenzhen exceeds 40%, significantly higher than that obtained in Shijiazhuang. Such high reduction in commuting distance of over 40% may be due to the large disparity in housing price in Shenzhen, creating a barriers for households to choose the most convenient residence locations, leading to an extensive commuting by a large set of commuters. The result of adding new activity centers in the Shenzhen simulation is consistent with the results of Shijiazhuang, as shown in Figure S19B.
Supplementary Figure S19. $\Delta d$ of Shenzhen after SDGHS. (A) The reduction in commuting distance after SDGHS with a different tolerances. (B) The reduction in commuting distance after SDGHS with a medium tolerance and a different number of new activity centers.
Home Swapping in Abidjan. We utilize data from Abidjan, the most economically prosperous center in Cote d’Ivoire, for the simulation of home swapping. First, based on the location of commuters’ residence and workplace, we visualize the spatial distributions of residence and workplace in Figure S20, where the transition from orange to deep red indicates areas with a higher concentration of residences and workplaces, suggesting a high population density. Conversely, green and blue areas represent locations with fewer residences and workplaces. By computing the KL divergence between the two spatial distributions, we found that the KL divergence is 0.113, suggesting distributions of residence and workplace in Abidjan are highly similar.
Supplementary Figure S20. The spatial distributions of residence and workplace. (A) shows the heat map of residence distribution in Abidjan; (B) shows the heat map of workplace distribution in Abidjan.
The distributions of residence and workplace in Abidjan are similar, implying a significant room for exploitation using GHS. The distribution of individual commuting distance of commuters in Abidjan is shown in Figure S21. As we can see, most commuters travel 5 km to their work, and there are a large number of commuters who travel for a commuting distance exceeds 10 km.
Supplementary Figure S21. The distribution of commuting distance in Abidjan.
After simulation of GHS in Abidjan, we found that single-commuter household achieved an 83.66% reduction in commuting distance. Since we cannot infer household membership from this simple data, we used the same methodto form double-commuter householdsby randomly pairing individuals living at the same latitude and longitude, resulting in a reduction of 41.36% of commuting distance after GHS as shown in Figure S22A.
Supplementary Figure S22. $\Delta d$ after GHS. (A) shows the reduction in commuting distance after GHS for households of different size; (B) shows $\Delta d$ as a function of the ratio of single-commuter to double-commuter households.
Next, we examine the reduction in commuting distance as a function of households of different size in Figure S22A. As we can see, $\Delta d$ significantly decreases from single-commuter to double-commuter households, while it decreases slowly for households with three, four, or five commuters. The highest $\Delta d$ is found in single-commuter households, and both single and double-commuter households can represent the overall benefit brought by GHS to the city. It is because the city is not composed purely of single or double-commuter families, but rather a certain proportion of single and double-commuter households. Figure S22B shows $\Delta d$ after GHS as a function of the ratio of single-commuter to double-commuter households, showing that as the proportion of single-commuter households increases, $\Delta d$ gradually increases. This suggests that, in comparison to small- or medium-sized cities, metropolitan areas with a higher percentage of single-commuter households have a higher potential for reducing commuting distance by GHS.
Supplementary Figure S23. $\Delta d$ of Abidjan after c-GHS. (A) $\Delta d$ after c-GHS. (B) $\Delta d$ after c-GHS with $T_{center}=3km$ for different number of new activity centers.
From Figure S23A, it can be observed that after c-GHS, $\Delta d$ improves as the restrictions are relaxed. Then we examined the effect of adding new activity centers, which are established by the most-visited locations as new activity centers at a distance of 3km from the original center. In Figure S23B, the result after c-GHS with new activity centers in Abidjan is consistent with the results of Shijiazhuang in the main text.